OpenStudy (anonymous):

A floor 5m by 20m is covered by square tiles of side 20cm. How many tiles are needed ?

OpenStudy (anonymous):

Any one ?

OpenStudy (anonymous):

area of floor= 5*20= 100 m area of tile = 20cm*20cm= 400 cm= 4m tiles required= 100/4= 25

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

whats the correct answer??

OpenStudy (anonymous):

2500 tiles

OpenStudy (anonymous):

it is 5 times 5 times 5 times 20

OpenStudy (anonymous):

yes, i am sorry.. u will divide 400 cm by 10000 to convert it into meters i.e. 400/10000=0.04m then divide 100 by 0.04 and u'll get 2500.

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

any way thanks